OpenStudy (anonymous):

Simplify: (4xy^-2) / [ 12x^(-1/3) y^5] I'll write the equation nicer w/ the equation box

OpenStudy (anonymous):

\[(4xy ^{-2}) / (12x ^{-1/3}y ^{-5})\]

OpenStudy (chaise):

4xy^-2 = 4/xy^2 12x^-(1/3) = 12/cubroot(12x) y^-5 = 1/y^5 Can you do the rest?

OpenStudy (anonymous):

lemem try

OpenStudy (anonymous):

wait what am i supposed to do after your step?

OpenStudy (anonymous):

Chaise, I think it is \[\left(\begin{matrix}4x y^{-2} \\ 12x ^{-1/3}y ^{-5}\end{matrix}\right) = \left(\begin{matrix}4x ^{1}x ^{1/3}y ^{5} \\ 12y ^{2}\end{matrix}\right)\] The x doesn't go to the denominator, right?

OpenStudy (anonymous):

so whose correct??

OpenStudy (anonymous):

@denebel is this ur final answer

OpenStudy (anonymous):

i dont think atht is fully simplified

OpenStudy (anonymous):

No that is not the final answer. You need to simplify it some more. For instance, with the x, there are 2 x's: one to the first power, and the other to the 1/3 power. Do you remember what you can do with them?

OpenStudy (anonymous):

ur supposed to add them right?

OpenStudy (anonymous):

can you show out teh final solution and the steps?

OpenStudy (anonymous):

Yes, you are supposed to add them. You are almost there... Then you can simplify the 4/12 right?

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

Okay... now put it all together.

OpenStudy (anonymous):

I got : \[(x ^{4/3}y ^{5}) / (12y ^{2})\] is taht correct

OpenStudy (anonymous):

Why is the 12 there?

OpenStudy (anonymous):

oh i meant 3y^2

OpenStudy (anonymous):

That is correct.

OpenStudy (anonymous):

is teh rest correct? btw do we not have to do anything to the x^(4/3)? is taht ok to have it liek taht

OpenStudy (anonymous):

\[(x ^{4/3}y ^{3})/3\]

OpenStudy (anonymous):

Oh, yes, Arnab09 is correct, we forgot about the y.

OpenStudy (dumbcow):

Basically you are just subtracting the exponents when you divide variables.